Car A and car B are travelling at 7 kph and 21 kph, respectively. If they are travelling on perpendicular roads, after how many hours will they be 75 kilometers apart?
Answer by stanbon(75887)

Car A and car B are travelling at 7 kph and 21 kph, respectively. If they are travelling on perpendicular roads, after how many hours will they be 75 kilometers apart?
---------
Car A distance for x hrs:: 7x
Car B distance for x hrs:: 21x
-------
Equation:
(7x)^2 + (21x)^2 = 75^2
(49 + 441)x^2 = 5625
-------
490x^2 = 5625
x^2 = 11.48
x = 3.39 hrs
------
Cheers,
Stan H.
-------------------